# Envelopes in the class of Banach algebras of polynomial growth and $C^{\infty}$-functions of a finite number of free variables

### v4

@@ Abstract.
We introduce the notion of envelope of a topological algebra (in particular, an arbitrary associative algebra) with respect to a class of Banach algebras. In the case of the class of real Banach algebras of polynomial growth, i.e., admitting a $C^{\infty}$-functional calculus for every element, we get a functor that maps the algebra of polynomials in $k$ variables to the algebra of $C^{\infty}$-functions on $\mathbb{R}^{k}$. The envelope of a general commutative or non-commutative algebra can be treated as an algebra of $C^{\infty}$-functions on some commutative or non-commutative space. In particular, we describe the envelopes of the universal enveloping algebra of finite-dimensional Lie algebras, the coordinate algebras of the quantum plane and quantum $SL(2)$ and also look at some commutative examples. A result on algebras of ‘free $C^{\infty}$-functions’, i.e., the envelopes of free associative algebras of finite rank $k$, is announced for general $k$ and proved for $k\leqslant 2$.
@@
Banach algebra of polynomial growth, envelope with respect to a class of Banach algebras, free $C^{\infty}$-function, universal enveloping algebra, quantum plane, quantum SL(2)
@@
Here we consider a construction which, at first sight, should be even more limited in its possibilities than the Arens–Michael envelope but, in fact, its application to certain classical non-commutative algebras demonstrates the non-triviality of the theory and links it to various areas of algebra and analysis. Namely, we propose to use the envelope with respect to the class of Banach algebras of polynomial growth as a bridge to the $C^{\infty}$-case. A real unital Banach algebra $B$ is said to be of polynomial growth if for every $b\in B$ there are $K>0$ and $\alpha\geqslant 0$ such that $\|e^{isb}\|\leqslant K(1+|s|)^{\alpha}$ for all $s\in\mathbb{R}$; equivalently, $b$ admits a $C^{\infty}$-functional calculus; see [5]. (Note that the theory only makes sense if the ground field is $\mathbb{R}$).
@@
Although theoretical constructions occupy a considerable part of this text, the emphasis is on examples. In addition to free algebras, we consider the universal enveloping algebras of finite-dimensional Lie algebras (refining results in [5]), the coordinate algebras of the quantum plane and quantum group $SL_{q}(2,\mathbb{R})$; see § 4. The commutative case is also of interest and discussed in § 3. We discuss the general properties of the class $\mathsf{PGL}$ in § 2 and also give a universal construction of $C^{\infty}$-tensor algebra in § 1. Note that the $C^{\infty}$-tensor algebra of a $k$-dimensional space is exactly the algebra of ‘free $C^{\infty}$-functions’ of rank $k$.
@@
Note that (in the case of real algebras) the second and third classes both contain $\mathsf{PG}$, which plays major role in this article. The enveloping functor with respect to the third class is introduced in [7] and also some examples are given there. The stability properties for strictly real Banach algebras are proved in [31, Theorem 4.4] but the envelopes have not been studied yet. Note also that for algebras with involution, envelopes with respect to the class of $C^{*}$-algebras have been studied earlier; see, for example, [19, 27].
@@ 1. Local containment and envelopes
We denote by $\mathsf{TA}$ the category of topological algebras over $\mathbb{C}$ or $\mathbb{R}$, that is, topological vector spaces endowed with separately continuous multiplication; see [30, Chapter 1, § 1, p. 4, Definition 1.1] or [20, § 1, p. 6, Definition 1.6]. In general, the underlying vector space is not assumed to be locally convex or complete and the existence of identity is also not assumed. Here the morphisms are continuous homomorphisms. Note that a subclass of objects in $\mathsf{TA}$ can be treated as a full subcategory. Following the tradition of functional analysis, we use the somewhat old-fashioned term projective limit (aka ‘inverse limit’) for what is called directed limit in category theory.
@@
Recall that a topological algebra whose topology can be determined by a family of submultiplicative seminorms is called an Arens–Michael algebra. If $\mathsf{C}$ is contained in the class of all Banach algebras, then each algebra locally contained in $\mathsf{C}$ is obviously an Arens–Michael algebra. Furthermore, an Arens–Michael algebra is exactly an algebra that is locally in the class of all Banach algebras. Note also that the limit of a projective system of Banach algebras in $\mathsf{TA}$ coincides with the limit in the category of Arens–Michael algebras; see, e.g., [30, Chapter 3, § 2, p. 84, (2.8)].
@@
The choice of the term ‘envelope’ is more likely due to historical considerations; cf. [21, Chapter 5]. It could just as well be called ’free functor’. However, we reserve the last term for those cases where the forgetful functor changes the algebraic type of the category (such as ‘associative algebras $\rightarrow$ vector spaces’ or ‘associative algebras $\rightarrow$ sets’). Akbarov in his papers [2, 3] also uses the term ‘envelope’, which has a similar but different meaning.
@@
In [17], Dixon considered varieties of Banach algebras, which are defined as classes of Banach algebras closed under passing to closed subalgebras, quotient algebras, bounded products and images under isometric isomorphisms. The conditions we impose on a class of Banach algebras are weaker. In particular, it is easy to see that $\mathsf{PG}$ defined in § 2 and the class of Banach algebras satisfying a polynomial identity are not closed under bounded products and so are not varieties in the sense of Dixon.
@@
If $\mathsf{B}$ is the class of all Banach algebras, then the tensor $\mathsf{BL}$ algebra is called the analytic tensor algebra (or Arens-Michael tensor algebra) and denoted by $\widehat{T}(E)$. In the case of the ground field $\mathbb{C}$, see a proof of the existence and an explicit construction in [34, § 4.2]. For $\mathbb{R}$ the proof is identical.
@@ 2. Banach algebras of polynomial growth and $\mathsf{PGL}$ algebras
(A) An element $b$ of a real Banach algebra is of polynomial growth if it satisfies to the (equivalent) Conditions (1) and (2) of Lemma 2.1 (or (2.1) in the unital case).
@@
For unital algebras, the assumption in the definition is a special case of that in [5, Definition 2.6]. Note also that the term ‘real generalized scalar operator’ is sometimes used for an operator satisfying (2.1).
@@
Let $p\in[1,\infty)$. We claim that the real Banach algebra $\ell_{p}$ (with the pointwise multiplication) is of polynomial growth. It suffices to check that $s\mapsto\|e^{isb}-1\|$ has polynomial growth for every $b\in\ell_{p}$. Take $b=(x_{n})\in\ell_{p}$. Since $e^{isb}-1=(e^{isx_{n}}-1)$, we have
@@
The following structural result is useful. Here we denote by $\mathop{\mathrm{Rad}}\nolimits B$ the Jacobson radical of an associative algebra $B$.
@@
[5, Theorem 2.8 and Proposition 2.9] Let $B$ be a unital Banach algebra of polynomial growth. Then $B/\mathop{\mathrm{Rad}}\nolimits B$ is commutative and $\mathop{\mathrm{Rad}}\nolimits B$ is nilpotent.
@@
3. Envelopes of certain commutative algebras
@@
The envelope of every locally convex algebra with respect to $\mathsf{PG}$ exists and is unique up to natural isomorphism.
@@
Although our main object of study are envelopes of non-commutative algebras, we first consider the commutative case. The following result is the main reason that led the author to study envelopes with respect to $\mathsf{PG}$.
@@
Let $b$ be an element of polynomial growth in a real Banach algebra $B$ and $V$ an open subset of $\mathbb{R}$ containing $\mathop{\mathrm{Sp}}\nolimits b$. Then there exists a unique multiplicative functional calculus, i.e., a continuous homomorphism $\Phi_{V}\!:C^{\infty}(V)\to B$ extending the homomorphism $\mathbb{R}[x]\to B\!:x\mapsto b$.
@@
Suppose that $B\in\mathsf{PG}$ and $b$ is an invertible element of $B$. Then $0\notin\mathop{\mathrm{Sp}}\nolimits b$ and thus $\mathbb{R}^{\times}$ is an open neighborhood of $\mathop{\mathrm{Sp}}\nolimits b$. Since $b$ is of polynomial growth, we can apply Proposition 3.5 with $V=\mathbb{R}^{\times}$. Hence there is a unique continuous homomorphism $C^{\infty}(\mathbb{R}^{\times})\to B$ extending $\mathbb{R}[x,x^{-1}]\to B$. Since $C^{\infty}(\mathbb{R}^{\times})\in\mathsf{PGL}$ by Corollary 2.8, this means that $\mathbb{R}[x,x^{-1}]\to C^{\infty}(\mathbb{R}^{\times})$ is an envelope with respect to $\mathsf{PG}$. ∎
@@
We finish this section with examples in which we explicitly describe the envelopes of some finitely generated commutative algebras. The following proposition is useful.
Proposition 3.6.
It is obvious that Part (A) follows from Part (B). For a proof of Part (B) see [39, Chapitre VI, p. 119, Corollaire 1.5]. See also a proof of Part (A) in [29, Théorème 4] and partial cases in [23, 28].
@@
Let $A=\mathbb{R}[x,y]/(xy)$, where $(xy)$ is the ideal generated by $xy$. We first claim that the homomorphism
$\rho\!:\mathbb{R}[x,y]\to\mathbb{R}[t]\times\mathbb{R}[t]\!:h\mapsto(x\mapsto h(x,0),\,y\mapsto h(0,y))$
induced an isomorphism
$\rho^{\prime}\!:A\to\mathcal{R}(\Omega).$
where $\mathcal{R}(\Omega)\!:=\{(f,g)\in\mathbb{R}[t]\times\mathbb{R}[t]\!:\,f(0)=g(0)\}$. (Here $\Omega$ stands for the set of characters, which coincides with the union of the coordinate axes but we do not need this fact here.)
Indeed, $\rho$ maps $(xy)$ to 00 and so $\rho^{\prime}$ is well defined. If $\rho(h)=0$, then $h$ is divisible by $xy$ and so it is in $(xy)$. Thus $\rho^{\prime}$ is injective. On the other hand, if $f,g\in\mathbb{R}[t]$ and $f(0)=g(0)$, then we have $\rho(h)=(f,g)$ when $h(x,y)=f(x)+g(y)-f(0)$. It follows that $\rho^{\prime}$ is surjective and hence bijective.
We now claim that $\widehat{A}^{\,\mathsf{PG}}$ is given by the same restriction, i.e., the continuous extension of $\rho$ to a homomorphism $C^{\infty}(\mathbb{R}^{2})\to C^{\infty}(\mathbb{R})\times C^{\infty}(\mathbb{R})$ induces a topological isomorphism $\widehat{A}^{\,\mathsf{PG}}\to C^{\infty}(\Omega)$, where
$C^{\infty}(\Omega):=\{(f,g)\in C^{\infty}(\mathbb{R})\times C^{\infty}(\mathbb{R})\!:\,f(0)=g(0)\}.$
Let $B$ be a Banach algebra of polynomial growth and $\varphi\!:A\to B$ be a homomorphism. We identify $A$ with a subalgebra of $\mathbb{R}[t]\times\mathbb{R}[t]$ as above and denote $(t,0)$ and $(0,t)$ by $X$ and $Y$, respectively. Let $\Psi_{X}$ and $\Psi_{Y}$ be the $C^{\infty}$-functional calculi $C^{\infty}(\mathbb{R})\to B$ associated the elements $\varphi(X)$ and $\varphi(Y)$; see Proposition 3.3. Then the formula
$C^{\infty}(\Omega)\to B\!:(f,g)\mapsto\Psi_{X}(f)+\Psi_{Y}(g)-f(0)$
determines a continuous linear map $\widehat{\varphi}$ that is a homomorphism since $XY=0$. Thus $\widehat{\varphi}$ is a desired extension of $\varphi$. Finally, it is easy to see that $\mathcal{R}(\Omega)$ is dense in $C^{\infty}(\Omega)$. Therefore $\widehat{\varphi}$ is unique.
4. Envelopes of certain non-commutative algebras
In general, the envelope of a commutative algebra with respect to $\mathsf{PG}$ can be treated as an algebra of $C^{\infty}$-functions on some commutative space and, moreover, the envelope of a non-commutative algebra as an algebra of $C^{\infty}$-functions on a non-commutative space. In this section, we consider first non-commutative examples: the envelopes of the coordinate algebras of quantum planes and quantum groups $SL_{q}(2,\mathbb{R})$ and also of the universal enveloping algebras of finite-dimensional Lie algebras. But we start with another example.
Example 4.1.
Consider the universal real algebra $A$ with generators $u$ and $v$ satisfying the relation $vu=1$. Its $C^{*}$-version is called the Toeplitz algebra or the ‘functional algebra on the quantum disc’; see, e.g., [16, 32]. Its ‘holomorphic’ [33, 36] and ‘smooth’ versions [14, 15] are also of interest. The latter is an extension of $C^{\infty}(S^{1})$ with respect to a certain ideal. Of course, it would be natural to expect the ‘smooth Toeplitz algebra’ coincides with the envelope of $A$ with respect to $\mathsf{PG}$. But this is not the case. In fact, the envelope satisfies rigid restrictions and we get $C^{\infty}(\mathbb{R}^{\times})$.
@@ 4. Envelopes of certain non-commutative algebras
Put $\mathcal{R}(\Omega)\!:=\{(f,g)\in\mathbb{R}[t]\times\mathbb{R}[t]\!:\,f(0)=g(0)\}$ and $C^{\infty}(\Omega)\!:=\{(f,g)\in C^{\infty}(\mathbb{R})\times C^{\infty}(\mathbb{R})\!:\,f(0)=g(0)\}$ as in Example 3.7. Denote $(t,0)$ and $(0,t)$ by $X$ and $Y$ and consider the linear maps
@@
In [5], to prove results of this kind the author used a direct approach with estimation of seminorms. Here an indirect approach is applied. It is based on Proposition 3.6 combined with the following simple lemma on products in the category of complete locally convex spaces. Recall that a continuous map is called is topologically injective if it is a homeomorphism on its range.
@@
  • If each $\varphi_{i}$ has closed range, then so is $\varphi$.
@@
(B) In the category of complete locally convex spaces, a morphism is an extremal monomorphism if and only if it a topologically injective map. So Part (B) follows from general properties of products and extremal monomorphisms; see, e.g., [1, Propositions 10.15(2) and 10.26(2)]. ∎
We need the following result on ordered calculus, which also be used in § 5
@@
[5, Theorem 3.3] Let $b_{1},\dots,b_{m}$ be elements of an Arens-Michael $\mathbb{R}$-algebra $B$. Suppose that $b_{1},\dots,b_{k}$ ($k\leqslant m$) are of polynomial growth and $b_{k+1},\dots,b_{m}$ are nilpotent. Then the linear map $\mathbb{R}[\lambda_{1},\dots,\lambda_{m}]\to B$ taking the (commutative) monomial $\lambda_{1}^{\beta_{1}}\cdots\lambda_{m}^{\beta_{m}}$ to the (non-commutative) monomial $b_{1}^{\beta_{1}}\cdots b_{m}^{\beta_{m}}$ extends to a continuous linear map
@@
(A) It follows from Corollary 2.8 that it suffices to construct a homomorphism from $\mathcal{R}(\mathbb{R}^{2}_{q})$ to a product of algebras of the form $C^{\infty}(M,\mathrm{T}_{p})$ and extend it to a topologically injective continuous linear map defined on $C^{\infty}(\Omega)\mathbin{\widehat{\otimes}}\mathbb{R}[[u]]$. This basic idea is the same as in the proof of Theorem 4.3 in [5] and we also use it in the proofs of Theorems 4.10 and 5.6.

### v5

@@ Abstract.
We introduce the notion of envelope of a topological algebra (in particular, an arbitrary associative algebra) with respect to a class of Banach algebras. In the case of the class of real Banach algebras of polynomial growth, i.e., admitting a $C^{\infty}$-functional calculus for every element, we get a functor that maps the algebra of polynomials in $k$ variables to the algebra of $C^{\infty}$-functions on $\mathbb{R}^{k}$. The envelope of a general commutative or non-commutative algebra can be treated as an algebra of $C^{\infty}$-functions on some commutative or non-commutative space. In particular, we describe the envelopes of the universal enveloping algebra of finite-dimensional Lie algebras, the coordinate algebras of the quantum plane and quantum group $SL(2)$ and also look at some commutative examples. A result on algebras of ‘free $C^{\infty}$-functions’, i.e., the envelopes of free associative algebras of finite rank $k$, is announced for general $k$ and proved for $k\leqslant 2$.
@@
Banach algebra of polynomial growth, envelope with respect to a class of Banach algebras, free $C^{\infty}$-function, universal enveloping algebra, quantum plane, quantum group SL(2)
@@
Here we consider a construction which, at first sight, should be even more limited in its possibilities than the Arens–Michael envelope but, in fact, its application to certain classical non-commutative algebras demonstrates the non-triviality of the theory and links it to various areas of algebra and analysis. Namely, we propose to use the envelope with respect to the class of Banach algebras of polynomial growth as a bridge to the $C^{\infty}$-case. A real unital Banach algebra $B$ is said to be of polynomial growth if for every $b\in B$ there are $K>0$ and $\alpha\geqslant 0$ such that $\|e^{isb}\|\leqslant K(1+|s|)^{\alpha}$ for all $s\in\mathbb{R}$; equivalently, $B$ admits a $C^{\infty}$-functional calculus associated with $b$; see [5]. (Note that the theory only makes sense if the ground field is $\mathbb{R}$).
@@
Although theoretical constructions occupy a considerable part of this text, the emphasis is on examples. In addition to free algebras, we consider the universal enveloping algebras of finite-dimensional Lie algebras (refining results in [5]), the coordinate algebras of the quantum plane and quantum group $SL_{q}(2,\mathbb{R})$; see § 3. The commutative case is also of interest and discussed there. We examine the general properties of the class $\mathsf{PGL}$ in § 2 and also give a universal construction of $C^{\infty}$-tensor algebra in § 1. Note that the $C^{\infty}$-tensor algebra of a $k$-dimensional space is exactly the algebra of ‘free $C^{\infty}$-functions’ of rank $k$.
@@
Note that (in the case of real algebras) the second and third classes both contain $\mathsf{PG}$, which plays major role in this article. The enveloping functor with respect to the third class is introduced in [7] and also some examples are given there. The stability properties for strictly real Banach algebras are proved in [32, Theorem 4.4] but the envelopes have not been studied yet. Note also that for algebras with involution, envelopes with respect to the class of $C^{*}$-algebras have been studied earlier; see, for example, [20, 28].
@@ 1. Local containment and envelopes
We denote by $\mathsf{TA}$ the category of topological algebras over $\mathbb{C}$ or $\mathbb{R}$, that is, topological vector spaces endowed with separately continuous multiplication; see [31, Chapter 1, § 1, p. 4, Definition 1.1] or [21, § 1, p. 6, Definition 1.6]. In general, the underlying vector space is not assumed to be locally convex or complete and the existence of identity is also not assumed. Here the morphisms are continuous homomorphisms. Note that a subclass of objects in $\mathsf{TA}$ can be treated as a full subcategory. Following the tradition of functional analysis, we use the somewhat old-fashioned term projective limit (also known as ‘inverse limit’) for what is called directed limit in category theory.
@@
Recall that a complete topological algebra whose topology can be determined by a family of submultiplicative seminorms is called an Arens–Michael algebra. If $\mathsf{C}$ is contained in the class of all Banach algebras, then each algebra locally contained in $\mathsf{C}$ is obviously an Arens–Michael algebra. Furthermore, an Arens–Michael algebra is exactly an algebra that is locally in the class of all Banach algebras. Note also that the limit of a projective system of Banach algebras in $\mathsf{TA}$ coincides with the limit in the category of Arens–Michael algebras; see, e.g., [31, Chapter 3, § 2, p. 84, (2.8)].
@@
The choice of the term ‘envelope’ is more likely due to historical considerations; cf. [22, Chapter 5]. It could just as well be called ’free functor’. However, we reserve the last term for those cases where the forgetful functor changes the algebraic type of the category (such as ‘associative algebras $\rightarrow$ vector spaces’ or ‘associative algebras $\rightarrow$ sets’). Akbarov in his papers [2, 3] also uses the term ‘envelope’, which has a similar but different meaning.
@@
In [18], Dixon considered varieties of Banach algebras, which are defined as classes of Banach algebras closed under passing to closed subalgebras, quotient algebras, bounded products and images under isometric isomorphisms. The conditions we impose on a class of Banach algebras are weaker. In particular, it is easy to see that $\mathsf{PG}$ defined in § 2 and the class of Banach algebras satisfying a polynomial identity are not closed under bounded products and so they do not form a variety in the sense of Dixon.
@@
If $\mathsf{B}$ is the class of all Banach algebras, then the tensor $\mathsf{BL}$ algebra is called the analytic tensor algebra (or Arens–Michael tensor algebra) and denoted by $\widehat{T}(E)$. In the case of the ground field $\mathbb{C}$, see a proof of the existence and an explicit construction in [35, § 4.2]. For $\mathbb{R}$ the proof is identical.
@@ 2. Banach algebras of polynomial growth and $\mathsf{PGL}$ algebras
(A) An element $b$ of a real Banach algebra is of polynomial growth if it satisfies the (equivalent) Conditions (1) and (2) of Lemma 2.1 (or (2.1) in the unital case).
@@
For unital algebras, the assumption in the preceding definition is a special case of that in [5, Definition 2.6]. Note also that the term ‘real generalized scalar operator’ is sometimes used for an operator satisfying (2.1).
@@
Let $p\in[1,\infty)$. We claim that the real Banach algebra $\ell_{p}$ (with pointwise multiplication) is of polynomial growth. It suffices to check that $s\mapsto\|e^{isb}-1\|$ has polynomial growth for every $b\in\ell_{p}$. Take $b=(x_{n})\in\ell_{p}$. Since $e^{isb}-1=(e^{isx_{n}}-1)$, we have
@@
The following structural result is useful. Here we denote by $\operatorname{Rad}B$ the Jacobson radical of an associative algebra $B$.
@@
[5, Theorem 2.8 and Proposition 2.9] Let $B$ be a unital Banach algebra of polynomial growth. Then $B/\operatorname{Rad}B$ is commutative and $\operatorname{Rad}B$ is nilpotent.
@@
3. Envelopes of certain commutative and non-commutative algebras
In general, the envelope of a commutative algebra with respect to $\mathsf{PG}$ can be treated as an algebra of $C^{\infty}$-functions on some commutative space and, moreover, the envelope of a non-commutative algebra as an algebra of $C^{\infty}$-functions on a non-commutative space. In this section, we consider first the commutative case and next the following non-commutative examples: the envelopes of the coordinate algebras of quantum planes, the quantum groups $SL_{q}(2,\mathbb{R})$ and the universal enveloping algebras of finite-dimensional Lie algebras.
@@ 3. Envelopes of certain commutative and non-commutative algebras
The envelope of every locally convex algebra with respect to the class $\mathsf{PG}$ exists and is unique up to natural isomorphism.
@@
The commutative case
Although our main object of study are envelopes of non-commutative algebras, we first look at the commutative case. The following result is the main reason that led the author to study envelopes with respect to $\mathsf{PG}$.
@@
Let $b$ be an element of polynomial growth in a real Banach algebra $B$ and $V$ an open subset of $\mathbb{R}$ containing $\operatorname{Sp}b$. Then there exists a unique multiplicative functional calculus, i.e., a continuous homomorphism $\Phi_{V}\!:C^{\infty}(V)\to B$ extending the homomorphism $\mathbb{R}[x]\to B\!:x\mapsto b$.
@@
Suppose that $B\in\mathsf{PG}$ and $b$ is an invertible element of $B$. Then $0\notin\operatorname{Sp}b$ and thus $\mathbb{R}^{\times}$ is an open neighborhood of $\operatorname{Sp}b$. Since $b$ is of polynomial growth, we can apply Proposition 3.5 with $V=\mathbb{R}^{\times}$. Hence there is a unique continuous homomorphism $C^{\infty}(\mathbb{R}^{\times})\to B$ extending $\mathbb{R}[x,x^{-1}]\to B$. Since $C^{\infty}(\mathbb{R}^{\times})\in\mathsf{PGL}$ by Corollary 2.8, this means that $\mathbb{R}[x,x^{-1}]\to C^{\infty}(\mathbb{R}^{\times})$ is an envelope with respect to $\mathsf{PG}$. ∎
@@
First non-commutative example
We start our discussion of non-commutative algebras with a simple example.
@@
Consider the universal real algebra $A$ with generators $u$ and $v$ satisfying the relation $vu=1$. Its $C^{*}$-version is called the Toeplitz algebra or the ‘functional algebra on the quantum disc’; see, e.g., [17, 33]. Its ‘holomorphic’ [34, 37] and ‘smooth’ versions [14, 15] are also of interest. The latter is an extension of $C^{\infty}(S^{1})$ with respect to a certain ideal. Of course, it would be natural to expect that the ‘smooth Toeplitz algebra’ coincides with the envelope of $A$ with respect to $\mathsf{PG}$. But this is not the case. In fact, the envelope satisfies rigid restrictions and we get $C^{\infty}(\mathbb{R}^{\times})$.
@@
Put $\mathcal{R}(\Omega)\!:=\{(f,g)\in\mathbb{R}[t]\times\mathbb{R}[t]\!:\,f(0)=g(0)\}$ and $C^{\infty}(\Omega)\!:=\{(f,g)\in C^{\infty}(\mathbb{R})\times C^{\infty}(\mathbb{R})\!:\,f(0)=g(0)\}$. Denote $(t,0)$ and $(0,t)$ by $X$ and $Y$ and consider the linear maps
@@
In [5], to prove results of this kind the author used a direct approach with estimation of seminorms. Here an indirect approach is applied. It is based on a proposition on closed ideals and submodules combined with a simple lemma on products in the category of complete locally convex spaces.
Proposition 3.9.
It is obvious that Part (A) follows from Part (B). For a proof of Part (B) see [40, Chapitre VI, p. 119, Corollaire 1.5]. See also a proof of Part (A) in [30, Théorème 4] and partial cases in [24, 29].
Recall that a continuous map is called is topologically injective if it is a homeomorphism onto its range. We need a pair of lemmas on topological injectivity.
@@
  • If each $\varphi_{i}$ has closed range, then so does $\varphi$.
@@
(B) In the category of complete locally convex spaces, a morphism is an extremal monomorphism if and only if it is a topologically injective map. So Part (B) follows from general properties of products and extremal monomorphisms; see, e.g., [1, Propositions 10.15(2) and 10.26(2)]. ∎
Note that a family $(\psi_{ij}\!:X_{i}\to Y_{j};\,i\in I,\,j\in J)$ of continuous linear maps of complete locally convex spaces induces a continuous linear map $\prod_{i\in I}X_{i}\to\prod_{j\in J}Y_{j}$ under the assumption that for every $j$ the subset of $i$ such that $\psi_{ij}\neq 0$ is finite. Indeed, the finiteness condition implies that there is a morphism from $\prod_{i\in I}X_{i}$ to $Y_{j}$ for every $j$ and hence from $\prod_{i\in I}X_{i}$ to $\prod_{j\in J}Y_{j}$.
Lemma 3.11.
Let $(\psi_{nm}\!:X_{n}\to Y_{m};\,n,m\in\mathbb{N})$ be a family of continuous linear maps between Fréchet spaces such that $\psi_{nm}=0$ whenever $n>m$ (or whenever $n<m$) and $\psi_{mm}$ is topologically injective for every $m$. Then the map $\prod X_{n}\to\prod Y_{m}$ induced by this family is topologically injective.
Proof.
Consider the case when $\psi_{nm}=0$ for $n>m$. It follows from a theorem of Dugundji [19, Theorem 4.1] that a continuous linear map between Fréchet spaces is topologically injective if and only if it has a continuous left inverse map (possibly, non-linear).
is left inverse (cf. the formula for the inverse of a triangular $2\times 2$ matrix) and so we have the topological injectivity.
Further, it easily follows by induction that $\prod_{n\leqslant N}X_{n}\to\prod_{m\leqslant N}Y_{m}$ also has a continuous left inverse map for every $N\in\mathbb{N}$. Moreover, the sequence of left inverse maps is compatible with the projections $\prod^{N+1}X_{n}\to\prod^{N}X_{n}$ and $\prod^{N+1}Y_{m}\to\prod^{N}Y_{m}$. Since an infinite product is a projective limit of finite products (see, e.g., [1, Exercise 11B]), we have a continuous map $\prod Y_{m}\to\prod X_{n}$ that is, as easy to see, left inverse to $\prod X_{n}\to\prod Y_{m}$. Thus the latter map is topologically injective.
The argument for the case when $\psi_{nm}=0$ for $n<m$ is similar. ∎
We also need the following result on ordered calculus, which also will be used in § 4.
@@
[5, Theorem 3.3] Let $b_{1},\dots,b_{m}$ be elements of an Arens–Michael $\mathbb{R}$-algebra $B$. Suppose that $b_{1},\dots,b_{k}$ ($k\leqslant m$) are of polynomial growth and $b_{k+1},\dots,b_{m}$ are nilpotent. Then the linear map $\mathbb{R}[\lambda_{1},\dots,\lambda_{m}]\to B$ taking the (commutative) monomial $\lambda_{1}^{\beta_{1}}\cdots\lambda_{m}^{\beta_{m}}$ to the (non-commutative) monomial $b_{1}^{\beta_{1}}\cdots b_{m}^{\beta_{m}}$ extends to a continuous linear map
@@
(A) It follows from Corollary 2.8 that it suffices to construct a homomorphism from $\mathcal{R}(\mathbb{R}^{2}_{q})$ to a product of algebras of the form $C^{\infty}(M,{\mathrm{T}}_{p})$ and extend it to a topologically injective continuous linear map defined on $C^{\infty}(\Omega)\mathbin{\widehat{\otimes}}\mathbb{R}[[u]]$. This basic idea is the same as in the proof of Theorem 4.3 in [5] and we also use it in the proofs of Theorems 3.17 and 4.6.